OpenStudy (anonymous):

Find the vertex of the quadratic equation: y = 2x2 + 8x − 4

OpenStudy (anonymous):

Didn't know this one so I looked it up! 2(x^2 + 4x - 2) = y x^2 + 4x - 2 = y/2 x^2 + 4x +4 = y/2 + 6 (x+2)^2 = y/2 + 6 (x+2)^2 - 6 = y/2 2(x + 2)^2 - 12 = y so the vertex is (-2, -12)